OpenStudy (anonymous):

solve each system using substitution. Check your answer. 3x+2y=23 1/2x=4=y

Mehek (mehek14):

what does 1/2x = 4 = y mean?

OpenStudy (anonymous):

i meant subtract

Mehek (mehek14):

3x+2(1/2x-4)=23 3x + 1x - 8 = 23 4x - 8 = 23 4x = 31 x = 7.75

OpenStudy (anonymous):

thank you so much

Mehek (mehek14):

no problem